### v2

# Practical Experience with Stable Set and Coloring Relaxations

Abstract
Stable Set and Graph Coloring belong to the class of NP-hard optimization problems on graphs. It is well known that even near-optimal solutions for these problems are difficult to find in polynomial time. The Lovász theta function, introduced by Lovász in the late 1970s, provides a powerful tool in the study of these problems. It can be expressed as the optimal value of a semidefinite program and serves as a relaxation for both problems. Considerable effort has been devoted to investigating additional cutting planes to strengthen these relaxations. In our work, we use these models and consider new classes of cutting planes based on small cliques and cycles contained in the underlying graph. We demonstrate that identifying such violated constraints can be done efficiently and that they often lead to significant improvements over previous bounds. However, our computational experiments also show that the quality of these improvements may decrease with problem size, and in some instances, no improvement is observed.
Keywords: Lovász Theta Function, Stability Number, Chromatic Number, Semidefinite Programming
1 Introduction
We consider two fundamental combinatorial optimization problems, the stable set problem and the graph coloring problem. Given a graph $G$, a stable set is a subset of vertices such that no two vertices in the subset are adjacent. The cardinality of a maximum stable set is called the stability number of $G$ and is denoted by $\alpha(G)$. The complement of $G$ is denoted by $\overline{G}$. Thus, stable sets in $G$ are in one-to-one correspondence with complete subgraphs, also called cliques, in $\overline{G}$. The number $\omega(\overline{G}):=\alpha(G)$ denotes the cardinality of a maximum clique in $\overline{G}$.
A graph coloring is an assignment of colors to the vertices of a graph $G$ such that adjacent vertices receive distinct colors. This can be seen as a partitioning the vertices of the graph $G$ into $k$ stable sets $S_{1},\ldots,S_{k}$, since such a partition provides a $k$-coloring of $G$ by assigning the color $i$ to all vertices in $S_{i}$ for $i=1,\ldots,k$. The smallest number $k$ such that $G$ has a $k$-partition into stable sets is called the chromatic number of $G$ and is denoted by $\chi(G)$. Partitioning of the vertex set $V$ into complete subgraphs is called clique partitioning and the clique partition number is the smallest $k$ such that the graph has a clique partition into $k$ sets containing all edges. Therefore, the clique partition number of $\overline{G}$ coincides with the chromatic number of $G$. Thus, it is legitimate to consider only the stable set and the coloring problem, as done here.
The stable set and the coloring problem are both contained in Karp’s original list of NP-complete problems [23] from 1972. Hence, unless P = NP, no polynomial time algorithm can solve these problems exactly.
By definition, the inequality $\alpha(G)\leq\chi(\overline{G})$ holds for these NP-hard graph parameters. A well-known semidefinite relaxation for these parameters was introduced by Lovász in [27]. The Lovász theta function, which can be computed to a fixed precision in polynomial time (see, for instance, Vandenberghe and Boyd [44]), provides a bound that separates the stability number of a graph $G$ and the chromatic number of its complement $\overline{G}$:
$\alpha(G)\leq\vartheta(G)\leq\chi(\overline{G}).$ (1)
Since its introduction in 1979, the Lovász theta function has been extensively studied. Several authors considered various strengthenings that yield tighter bounds on $\alpha(G)$ and $\chi(G)$. A standard way to strengthen a relaxation is to add cutting planes; see, for instance, Fortet [11] and Barahona, Jünger, and Reinelt [3]. This approach has been exploited, for instance, in strengthenings proposed by Schrijver [38], Lovász and Schrijver [28], Szegedy [41], and Meurdesoif [29].
Another way to strengthen the Lovász theta function is to use a hierarchical approach. This systematic procedure involves strengthening a relaxation by adding additional variables and constraints, and it usually unfolds in multiple levels, each providing a tighter relaxation than the previous one. The most prominent hierarchies based on semidefinite programming (SDP) are the ones from Sherali and Adams [39], Lovász and Schrijver [28] and Lasserre [25]. Laurent investigated the computational application of these hierarchies to the stable set polytope in [26].
However, any strengthening of the Lovász theta function inevitably increases computational complexity. To address this, our research focuses on identifying subgraphs and structures where the separation problem can be efficiently solved and where the addition of valid inequalities into the SDP for computing the Lovász theta function can be executed within relatively short running times. In particular, in [35], we derived several inequalities valid for specific subgraphs containing maximal cliques. Our computational results indicated that incorporating these inequalities into the SDP formulation for the computation of the Lovász theta function significantly improves the bounds on both $\alpha(G)$ and $\chi(G)$.
Building on these promising results, we now investigate subgraphs induced either partially or entirely by odd cycles, and we derive several new valid inequalities for the stable set and graph coloring problems. Additionally, we conduct extensive computational experiments to demonstrate the effectiveness of these new constraints, alongside those we introduced in [35], in comparison to the previous state of the art.
The rest of this paper is organized as follows. In Section 2, we review the literature relevant for our work. Then, in Section 3 we present new inequalities valid for the stable set and graph coloring problems. Section 4 is devoted to computational experiments. Finally, we conclude with a short discussion of the results in Section 5.
We close this section with some words on notation used terminology. The vector of all-ones is denoted by $e$ and we write $J=ee^{T}$ for the matrix of all-ones. Furthermore, throughout this paper, $G=(V(G),E(G))$ denotes a simple undirected graph with $|V(G)|=n$ vertices and $|E(G)|=m$ edges. A graph $G^{\prime}$ is a subgraph of $G$ if $V(G^{\prime})\subseteq V(G)$ and $E(G^{\prime})\subseteq E(G)$. Let $V^{\prime}\subseteq V(G)$. We say that the subgraph $G^{\prime}$ of $G$ is induced by $V^{\prime}$ if $E(G^{\prime})$ consists of all edges $[i,j]$ with $i,j\in V^{\prime}$. A cycle $C$ in a graph $G$ is a connected subgraph of $G$ where all vertices have degree two. A cycle $C$ is an induced cycle in $G$ if the subgraph induced by $V(C)$ consists only of the edges in $C$. The length of a cycle is the number of its vertices.
2 Preliminaries
This section outlines the necessary preliminaries for our work. In Subsection 2.1, we formulate the stable set and graph coloring problems as integer programs and demonstrate how to derive semidefinite programming relaxations for these problems. Specifically, we show that the Lovász theta function provides both upper bounds on $\alpha(G)$ and lower bounds on $\chi(\overline{G})$. Subsection 2.2 then reviews existing enhancements of the Lovász theta function towards $\alpha(G)$ and $\chi(G)$, and outlines the primary objectives of this work in that context.
2.1 Stable Sets, Colorings, and the Lovász Theta Function
Both the stable set and the coloring problem can be formulated as integer optimization problems. The standard formulation for the stable set problem goes back to Padberg [33] and is given as follows. Let $x_{i}$ be a binary variable indicating whether the vertex $i$ is contained in a stable set ($x_{i}=1$) or not ($x_{i}=0$). Then, the stability number of a graph is the optimal value of the integer linear program
$\alpha(G)~{}=~{}\max\{e^{T}x\colon x_{i}+x_{j}\leq 1~{}\forall[i,j]\in E(G),~{}x_{i}\in\{0,1\}~{}\forall i\in V(G)\}.$ (2)
An alternative formulation is obtained by introducing quadratic constraints; see, for instance, Ben-Tal and Nemirovski [5]. Let $x_{i}$ be a binary variable as in formulation (2). Then, the binary constraint can be replaced by $x_{i}^{2}=x_{i}$, and the edge inequalities $x_{i}+x_{j}\leq 1$ can be written as $x_{i}x_{j}=0$. Thus,
$\alpha(G)~{}=~{}\max\{e^{T}x\colon x_{i}x_{j}=0~{}\forall[i,j]\in E(G),~{}x_{i}^{2}=x_{i}~{}\forall i\in V(G)\}.$ (3)
Since the computation of $\alpha(G)$ is NP-hard, one usually considers a suitable relaxation. A simple linear programming relaxation of the formulation (2) can be obtained by omitting the binary constraint and allowing
$0\leq x_{i}\leq 1~{}~{}\forall i\in V(G).$ (4)
However, as Nemhauser and Trotter noted in [31], this linear relaxation is relatively weak. To strengthen it, one can add inequalities that are valid for the convex hull of the set of incidence vectors of stable sets in $G$. We denote the set of these vectors as
$\mathcal{S}(G):=\{s\in\{0,1\}^{n}:s\mbox{ incidence vector of some stable set in }G\}.$
The stable set polytope $\operatorname{STAB}(G)$ is then defined as
$\operatorname{STAB}(G):=\operatorname{conv}\{s\colon s\in\mathcal{S}(G)\}.$
The study of valid inequalities for $\operatorname{STAB}(G)$ was initiated by Padberg in [33]. For instance, if $C$ denotes the vertex set of an odd cycle in $G$, then clearly at most $\frac{1}{2}(|V(C)|-1)$ vertices of $C$ can be contained in any stable set of $G$. Similarly, having a clique $Q$ in $G$, at most one of its vertices may be contained in any stable set. Thus, as shown in [33], the following two classes of inequalities are valid for $\operatorname{STAB}(G)$
[…]
In [19], Grötschel, Lovász, and Schrijver demonstrated that any linear function subject to the odd cycle constraints (5) can be optimized in polynomial time. However, optimizing a linear function with the clique constraints (6) is NP-hard; see, for instance, Nemhauser and Wolsey [32].
Besides the mentioned odd cycle and clique inequalities, many other inequalities valid for $\operatorname{STAB}(G)$ are known. For the sake of brevity, we list only some of these classes: odd antihole and odd wheel inequalities [33], web and antiweb inequalities introduced by Trotter in [43], rank inequalities introduced by Nemhauser and Trotter in [30]; some further inequalities can be found in works from Cheng and Cunningham [8], Cánovas, Landete, and Marín [6], and Cheng and de Vries [9].
Another way to obtain upper bounds on $\alpha(G)$ is to consider the Lovász theta function $\vartheta(G)$, introduced by Lovász [27] in his seminal paper from 1979. A semidefinite program to compute the Lovász theta function can be formulated in several ways, as demonstrated, for instance, by Grötschel, Lovász, and Schrijver in [18] and by Knuth in [24]. Here, we state the formulation that was given by Lovász and Schrijver in [28] and that can be directly derived by considering the integer programming formulation with quadratic constraints (3), as outlined, for instance, in Rendl [36].
Suppose that a vector $x\in\{0,1\}^{n}$ is feasible for the formulation (3) and consider the matrix $X=xx^{T}\in\{0,1\}^{n\times n}$. By definition, $X$ is positive semidefinite, and since $x$ is a binary vector, $x_{i}^{2}=x_{i}$ holds for every $i\in V(G)$. Hence, the main diagonal of $X$ is equal to $x$. Furthermore, for every edge, the entry $X_{i,j}$ equals zero. Nevertheless, the constraint $X=xx^{T}$ is not convex. Relaxing $X-xx^{T}=0$ to $X-xx^{T}\succeq 0$ and using the Schur complement
$X-xx^{T}\succeq 0~{}\Leftrightarrow~{}\begin{pmatrix}1&x^{t}\\ x&X\end{pmatrix}\succeq 0,$
we obtain the following SDP to compute the Lovász theta function
$\vartheta(G)~{}=~{}\max\Biggl{\{}e^{T}x\colon\begin{pmatrix}1&x^{T}\\ x&X\end{pmatrix}\succeq 0,~{}\mathrm{diag}(X)=x,~{}X_{i,j}=0~{}\forall[i,j]\in E(G)\Biggr{\}}.$ (7)
In [18], Grötschel, Lovász, and Schrijver introduced the theta body of a graph $G$ as the convex body $\operatorname{TH}(G)\subseteq\mathbb{R}^{n}$ given by the feasible region of the SDP to compute the Lovász theta function. Hence, considering the formulation (7), we obtain
$\operatorname{TH}(G):=\Biggl{\{}x\in\mathbb{R}^{n}~{}:~{}\exists X\in\mathbb{S}^{n}~{}:~{}\begin{pmatrix}1&x^{T}\\ x&X\end{pmatrix}\succeq 0,~{}\mathrm{diag}(X)=x,~{}X_{i,j}=0~{}\forall[i,j]\in E(G)\Biggr{\}}.$
The theta body has the advantage that we can optimize a linear function over $\operatorname{TH}(G)$ in polynomial time with an arbitrary fixed precision using semidefinite programming. Furthermore, by definition, we can see that the theta body contains the stable set polytope of $G$, i.e.,
$\operatorname{STAB}(G)\subseteq\operatorname{TH}(G)\subseteq\mathbb{R}^{n}.$
Additionally, feasibility for the problem in (7) also implies that $0\leq x_{i}\leq 1$. Moreover, Grötschel, Lovász, and Schrijver [18] showed that $\operatorname{TH}(G)$ satisfies all clique constraints. Therefore, $\vartheta(G)$ often provides a tight upper bound on the stability number $\alpha(G)$. This bound is generally stronger than those obtained from linear relaxations, as shown, for instance, by Balas, Ceria, Cornuéjols, and Pataki in [2].
In the next subsection, we will discuss the ways to strengthen the Lovász theta function $\vartheta(G)$ as an upper bound on $\alpha(G)$ by adding valid inequalities into the SDP for its computation. Clearly, inequalities valid for $\operatorname{STAB}(G)$ can be added into (7) in order to obtain better bounds. Nevertheless, we will also consider some inequalities valid for the matrix sets associated with $\operatorname{STAB}(G)$ and $\operatorname{TH}(G)$. Therefore, it will turn out to be useful to work with the matrix version of $\mathcal{S}(G)$. We denote it by $\mathcal{S}^{2}(G)$ and define it as
$\mathcal{S}^{2}(G):=\Biggl{\{}\begin{pmatrix}1\\ s\end{pmatrix}\begin{pmatrix}1\\ s\end{pmatrix}^{T}:s\in\operatorname{STAB}(G)\Biggr{\}}.$
Its convex hull is
$\operatorname{STAB}^{2}(G):=\operatorname{conv}\{\mathcal{S}^{2}(G)\}.$
Furthermore, we introduce the matrix set
$\mathcal{M}(G):=\Biggl{\{}X\in\mathbb{S}^{n}~{}:~{}\begin{pmatrix}1&x^{T}\\ x&X\end{pmatrix}\succeq 0,~{}\mathrm{diag}(X)=x,~{}X_{i,j}=0~{}\forall[i,j]\in E(G)\Biggr{\}}.$
Then, the projection of $\mathcal{M}(G)$ onto $\mathbb{R}^{n}$ is the theta body $\operatorname{TH}(G)$.
We now consider the graph coloring problem and show that the Lovász theta function also yields a lower bound on $\chi(G)$. To this end, we follow the presentation by Dukanovic and Rendl from [10]. Recall that a $k$-coloring of a graph G on $n$ vertices is a $k$-partition $(V_{1},\ldots,V_{k})$ of the vertex set $V(G)$ such that each $V_{i}$ is a stable set in $G$. We encode the $k$-partition by the characteristic vectors $s_{i}$ for $V_{i}$. Hence, $s_{i}\in\{0,1\}^{n}$ and for a $v\in V(G)$ we have $(s_{i})_{v}=1$ if $v\in V_{i}$, and 00 otherwise. Furthermore, the partition property implies that $s_{i}\neq 0$ and $\sum_{i=1}^{k}s_{i}=e$. The matrix
$X=\sum_{i=1}^{k}s_{i}s_{i}^{T}$
is called a coloring matrix. The convex hull of the set of all coloring matrices of $G$ is denoted by
$\operatorname{COL}(G):=\operatorname{conv}\{X:X\text{ is a coloring matrix of $G$}\}.$
Hence, a coloring matrix $X$ is the partition matrix associated to the $k$-partition $(V_{1},\ldots,V_{k})$. Clearly, $\operatorname{rank}(X)=k$ and $X$ is a symmetric $0/1$ matrix. Partition matrices can be characterized in several ways. In particular, a symmetric matrix $X\in\{0,1\}^{n\times n}$ is a $k$-partition matrix if and only if $\operatorname{diag}(X)=e$ and $tX-J\succeq 0$ for all $t\geq k$. Consequently, the chromatic number of a graph $G$ is the optimal value of the following SDP in binary variables
$\chi(G)~{}=~{}\min\{t\colon tX-J\succeq 0,~{}\mathrm{diag}(X)=e,~{}X_{i,j}=0~{}\forall[i,j]\in E(G),~{}X\in\{0,1\}^{n\times n}\}.$ (8)
Nevertheless, solving (8) is NP-hard. A tractable relaxation can be obtained by leaving out the binary condition. Furthermore,
$tX-J\succeq 0~{}\Leftrightarrow~{}\begin{pmatrix}t&e^{T}\\ e&X\end{pmatrix}\succeq 0$
by the Schur complement. Hence, we obtain the following semidefinite relaxation that yields a lower bound on the chromatic number $\chi(G)$
$\vartheta(\overline{G})~{}=~{}\min\Biggl{\{}t\colon\begin{pmatrix}t&e^{T}\\ e&X\end{pmatrix}\succeq 0,~{}\mathrm{diag}(X)=e,~{}X_{i,j}=0~{}\forall[i,j]\in E(G)\Biggr{\}}.$ (9)
Note that the optimal value of (9) is indeed the Lovász theta function $\vartheta(\overline{G})$. This can be demonstrated by considering the dual of the SDP problem (7) applied to the complement of $G$. In particular, this proves the Lovász sandwich theorem, as stated in equation (1) in the introduction of this paper.
2.2 Strengthening of the Lovász Theta Function
In the last subsection we have seen that the parameter $\vartheta(G)$ as an upper bound on $\alpha(G)$ is given as the optimal value for the semidefinite optimization problem (7). There are some obvious strengthenings towards $\alpha(G)$. The first strengthening $\vartheta^{\prime}(G)$ of the Lovász theta function was given by Schrijver [38] by adding the nonnegativity constraints
$X\geq 0.$ (10)
Gruber and Rendl [20] proposed a further strengthening $\vartheta^{\prime\Delta}(G)$ which is obtained by including the triangle inequalities
$\ X_{i,k}+X_{j,k}\leq X_{i,j}+x_{k},$$\quad\forall\{i,j,k\}\subseteq V(G)$ (11)
$x_{i}+x_{j}+x_{k}\leq 1+X_{i,k}+X_{j,k}+X_{i,j},$$\quad\forall\{i,j,k\}\subseteq V(G)$ (12)
into the SDP (7), yielding a chain of relaxations
$\alpha(G)\leq\vartheta^{\prime\Delta}(G)\leq\vartheta^{\prime}(G)\leq\vartheta(G).$
The validity of the triangle inequalities (11) and (12) for the stable set polytope was first observed by Padberg in [34]. In that work, Padberg introduced triangle inequalities valid for the boolean quadric polytope, which is the convex hull of all matrices $X$ of the form $xx^{T}$ where $x\in\{0,1\}^{n}$. Nevertheless, the polytope $\operatorname{STAB}(G)$ is derived by taking the face of the boolean quadric polytope defined by the equations $X_{i,j}=0$ for all $[i,j]\in E(G)$ and then projecting it onto the non-quadratic space. Consequently, the triangle inequalities valid for the boolean quadric polytope also hold for the stable set polytope.
Additionally, it is important to mention the relaxation proposed by Lovász and Schrijver in [28] which includes both the non-negativity constraint (10) and a subset of the triangle inequalities (11) and (12). This relaxation is derived using the Lift-and-Project operator introduced in the same work. While it does not encompass all triangle inequalities, making it slightly weaker than $\vartheta^{\prime\Delta}(G)$, it does satisfy all odd cycle inequalities (5). Therefore, incorporating the non-negativity constraint (10) along with the triangle inequalities (11) and (12) into the SDP (7) ensures that all odd cycle inequalities (5) are satisfied.
Several strengthenings of $\vartheta(\overline{G})$ towards $\chi(G)$ have been proposed and are quite similar to the strengthening of $\vartheta(G)$ towards $\alpha(G)$. Szegedy [41] introduced the nonnegativity constraint
$X\geq 0$ (13)
leading to $\vartheta^{+}(\overline{G})$. Meurdesoif [29] introduced an even tighter bound $\vartheta^{{+\Delta}}(\overline{G})$, which is obtained by adding the triangle inequalities
$X_{i,j}+X_{j,k}\leq X_{i,k}+1,\quad\forall\{i,j,k\}\subseteq V(G)$ (14)
into the SDP (9), resulting again in the chain of relaxations
$\vartheta(\overline{G})\leq\vartheta^{+}(\overline{G})\leq\vartheta^{+\Delta}(\overline{G})\leq\chi(G).$
We have just seen several rather obvious ways to tighten the Lovász theta function either towards $\alpha(G)$ or towards $\chi(G)$. In particular, any systematic tightening towards $\alpha(G)$ or $\chi(G)$ amounts to identifying a subset of linear inequalities which provide a reasonably accurate description of the underlying convex hull.
A general cutting plane theory along these lines is given by the class of Chvátal-Gomory cuts, see for instance Caprara and Fischetti [7]. Nevertheless, the identification of violated Chvátal-Gomory cuts is NP-hard in general. This motivates the search for problem-specific cutting planes which exploit the combinatorial structure of the underlying problem. Grötschel and Wakabayashi [17] consider the clique partition problem in $G$ which corresponds to coloring in $\overline{G}$ and describe several classes of cutting planes and investigate their practical impact, see [16]. Strengthening the relaxation for the stable set problem by adding odd cycle and triangle inequalities was studied by Gruber and Rendl in [20]. Computational experiments regarding strengthening both problems by adding triangle inequalities were done by Dukanovic and Rendl in [10] and Battista and De Santis in [4].
One of the most notable results in bounding the stability number $\alpha(G)$ and the chromatic number $\chi(G)$ based on the semidefinite programming relaxation was obtained recently by Gaar and Rendl in [12]. These bounds were achieved by employing the exact subgraph hierarchy introduced by Anjos, Adams, Rendl, and Wiegele in [1]. This hierarchy is particularly well-suited for classes of NP-hard problems based on graphs, where the projection of the problem onto a subgraph retains the same structure as the original problem.
In the context of the stable set and graph coloring problems, the authors proposed that for certain subsets of vertices, the corresponding submatrix should be contained in $\operatorname{STAB}^{2}(G)$ or $\operatorname{COL}(G)$ restricted to the respective subgraph. The exact subgraph hierarchy begins by computing the Lovász theta function at the first level of the hierarchy and then progressively considers all subsets of vertices with increasing cardinalities. While this method produces tight bounds on $\alpha(G)$ and $\chi(G)$, it is associated with significant computational challenges due to the complexity involved in describing the convex hull of the underlying subgraph.
Motivated by these challenges, our research presented in [35] focused on identifying certain subgraphs for which the description of $\operatorname{STAB}^{2}(G)$ and $\operatorname{COL}(G)$ is relatively simple and can be efficiently integrated into the SDP for computing the Lovász theta function. Specifically, we examined subgraphs induced by maximal cliques $Q\subseteq V(G)$ and $Q^{\prime}\subseteq V(G)$ and demonstrated that the following inequalities are valid for $\operatorname{STAB}^{2}(G)$
$\sum_{i\in V(Q)}X_{i,j}\leq X_{j,j},$ $\quad\forall Q\subseteq V(G),~{}\forall j\in V(G),~{}j\notin V(Q)$ (15)
$\sum_{i\in V(Q\cup Q^{\prime})}X_{i,i}\leq 1+\sum_{\begin{subarray}{c}i\in V(Q)\\ j\in V(Q^{\prime})\end{subarray}}X_{i,j},$ $\quad\forall Q\subseteq V(G),~{}Q^{\prime}\subseteq V(G),~{}Q\cap Q^{\prime}=\emptyset.$ (16)
Similarly, we established that the inequalities
$\sum_{i\in V(Q)}X_{i,k}$ $\leq 1,\quad\forall Q\subseteq V(G),~{}\forall k\in V,~{}k\notin V(Q)$ (17)
are valid for $\operatorname{COL}(G)$. Moreover, we showed that adding these inequalities along with the non-negativity constraints is sufficient to describe $\operatorname{STAB}^{2}(G)$ and $\operatorname{COL}(G)$ of respective subgraphs.
However, it is important to note that Giandomenico, Letchford, Rossi, and Smriglio arrived at a similar result while studying the strengthening of the linear relaxation for the stable set problem in [14]. Specifically, by considering the M(K,K) operator introduced by Lovász and Schrijver in [28], they derived the same class of inequalities (15) and (16). However, these inequalities were only applied to the linear relaxation. Further computational studies of these inequalities within the context of linear relaxation were conducted by Giandomenico, Rossi, and Smriglio in [15]. To our knowledge, our paper [35] is the first to explore the application of inequalities (15) and (16) in the SDP setting. Notably, our computational results from [35] demonstrate that incorporating these inequalities into the SDP to compute the Lovász theta function significantly improves bounds on both $\alpha(G)$ and $\chi(G)$.
This success has motivated our current study, where we explore additional inequalities valid for $\operatorname{STAB}^{2}(G)$ and $\operatorname{COL}(G)$ by focusing on specific subgraphs of $G$ that are easily separable. This time, these new inequalities are valid for subgraphs induced by odd cycles, either partially or entirely.
3 Some New Valid Inequalities
In this section, we present several new inequalities that are valid for $\operatorname{STAB}^{2}(G)$ and $\operatorname{COL}(G)$. We start by considering the stable set problem and present inequalities associated with odd cycles of length $5$.
Theorem 1.
Let $C$ denote a $5$-cycle in $G$. The following inequality is valid for $\operatorname{STAB}^{2}(G)$
$\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}X_{i,j}\leq 1.$ (18)
Proof.
Let $s=(s_{1},\ldots,s_{n})^{T}$ be an incidence vector of a stable set in $G$, and define $X=ss^{T}$. Then, we have
$\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}X_{i,j}=\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}s_{i}s_{j}.$
Since the stability number of a $5$-cycle is at most two, it follows that $|s\cap V(C)|\leq 2$, and therefore
$\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}s_{i}s_{j}\leq 1.$
Next, we explore the potential of strengthening the Lovász theta function $\vartheta(G)$ as an upper bound on $\alpha(G)$ by incorporating the inequalities (18) into the SDP formulation (7). Specifically, we first demonstrate that these newly introduced inequalities (18) imply the standard $5$-cycle inequalities (5). Importantly, the converse is not true: adding the standard $5$-cycle inequalities (5) does not imply the inequalities (18). Therefore, including constraints (18) into the SDP for the computation of $\vartheta(G)$ may yield better upper bounds on $\alpha(G)$ compared to using the Lovász theta function strengthened with the standard $5$-cycle inequalities (5).
In order to show these statements, we switch to the matrix set $\mathcal{M}(G)$. Its projection onto $\mathbb{R}^{n}$ is the theta body $\operatorname{TH}(G)$.
Lemma 2.
Let $C$ denote a $5$-cycle in $G$. If $X\in\mathcal{M}(G)$ and $\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}X_{i,j}\leq 1$, then
$\sum_{i\in V(C)}x_{i}\leq 2.$
Proof.
Let $X_{C}$ be the submatrix of $X$ induced by the cycle $C$. Since $X\in\mathcal{M}(G)$, the matrix $Y:=\begin{pmatrix}1&x^{T}_{c}\\ x_{c}&X_{c}\end{pmatrix}$, where $x_{c}=\operatorname{diag}(X_{c})$, is positive semidefinite. Hence, for any vector $a\in\mathbb{R}^{6}$ we have $a^{T}Ya\geq 0$. Consider the vector $a=(1,1,1,1,1,-2)^{T}$. Then, $a^{T}Ya\geq 0$ expands to
$2\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}X_{i,j}-3\sum_{i\in V(C)}x_{i}+4\geq 0,$
since $X_{i,j}=0$ for all $[i,j]\in E(C)$. Furthermore, since $\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}X_{i,j}\leq 1$, we can simplify this to
$3\sum_{i\in V(C)}x_{i}\leq 6,$
and therefore $\sum_{i\in V(C)}x_{i}\leq 2$. ∎
To demonstrate that the standard $5$-cycle inequality $\sum_{i\in V(C)}x_{i}\leq 2$ does not imply the inequality (7), we provide a counterexample. Let $C$ be an induced $5$-cycle with $V(C)=\{1,\dots,5\}$ and $E(C)=\{[1,2],[2,3],[3,4],[4,5],[1,5]\}$. Define the variables as follows
$X_{i,i}$ $:=0.4$ $\quad\forall i\in V(C)$
$x_{i}$ $:=X_{i,i}$ $\quad\forall i\in V(C)$
$X_{i,j}$ $:=0$ $\quad\forall[i,j]\in E(C)$
$X_{i,j}$ $:=0.209$ $\quad\forall[i,j]\notin E(C),$
Then, it is an easy exercise to show that $X\in\mathcal{M}(G)$. Furthermore, by construction, $\sum_{i\in V(C)}x_{i}=2$. However, we find that $\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}X_{i,j}=1.045>1$. Hence, this counterexample illustrates that satisfying $\sum_{i\in V(C)}x_{i}\leq 2$ does not necessarily imply the inequality (7).
The second inequality we introduce refers to an odd cycle $C$ of arbitrary length and a vertex not contained in $V(C)$.
Theorem 3.
Let $C$ be an odd cycle in $G$ and let $k\in V(G)$ such that $k\notin V(C)$. The following inequalities are valid for $\operatorname{STAB}^{2}(G)$
$\sum_{i\in V(C)}X_{i,k}$ $\leq\frac{1}{2}(|V(C)|-1)X_{k,k}$ (19)
$\sum_{i\in V(C)}X_{i,i}+\frac{1}{2}(|V(C)|-1)X_{k,k}$ $\leq\frac{1}{2}(|V(C)-1)+\sum_{i\in V(C)}X_{i,k}$ (20)
Proof.
We use the idea of the Lift-and-Project operator introduced in [28]. For any $k\in V(G)$ we have that $0\leq X_{k,k}\leq 1$. Hence, for a $k\notin V(G)$, multiplying the odd cycle inequality
$\sum_{i\in V(C)}X_{i,i}\leq\frac{1}{2}(|V(C)|-1)$
by $X_{k,k}$ and $1-X_{k,k}$ yields inequalities (19) and (20). ∎
In the context of the graph coloring problem, we introduce two new inequalities valid for the coloring polytope $\operatorname{COL}(G)$. Similar to the approach taken for the stable set problem, the first inequality pertains to odd cycles of length $5$, while the second refers to odd cycles of arbitrary length along with an additional vertex that is not contained in the cycle.
Theorem 4.
Let $C$ be a $5$-cycle in $G$. Then the following inequality is valid for $\operatorname{COL}(G)$
$\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}X_{i,j}\leq 2.$ (21)
Proof.
Let $X$ be a coloring matrix. Then $X=\sum_{p=1}^{\ell}s_{p}s_{p}^{T}$, where each $s_{p}$ is an incidence vector of a stable set in $G$ with $s_{p}\neq 0$ and $\sum_{p=1}^{\ell}s_{p}=e$. Hence, $\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}X_{i,j}$ counts the number of indices $p$ for which
$|s_{p}\cap V(C)|\geq 2.$ (22)
But the stability number of $C$ is at most two, and since $C$ is a $5$-cycle, there are at most two indices $p$ for which the inequality (22) holds. Hence, $\sum_{\begin{subarray}{c}i,j\in V(C)\\ i<j\end{subarray}}X_{i,j}\leq 2$. ∎
Theorem 5.
Let $C$ be an odd cycle in $G$ and let $k\in V(G)$ such that $k\notin V(C)$. Then the following inequality is valid for $\operatorname{COL}(G)$
$\sum_{i\in V(C)}X_{i,k}\leq\frac{1}{2}(|V(C)|-1).$ (23)
Proof.
Let $X$ be a coloring matrix, so $X=\sum_{p=1}^{\ell}s_{p}s_{p}^{T}$, where each $s_{p}$ is an incidence vector of a stable set in $G$ with $s_{p}\neq 0$ and $\sum_{p=1}^{\ell}s_{p}=e$. Therefore, the vertex $k$ is contained in a stable set $s_{p}$ for some $p\in\{1,\ldots,\ell\}$. Without loss of generality, assume $k\in s_{1}$. Then,
$\sum_{i\in V(C)}X_{i,k}=|s_{1}\cap V(C)|,$
and since the stability number of an odd cycle $C$ is at most $\frac{1}{2}(|V(C)|-1)$, we obtain the inequality (23). ∎
4 Computational Study
We now perform a computational study to evaluate the strengthening of the Lovász theta function $\vartheta(G)$ towards the stability number $\alpha(G)$ and chromatic number $\chi(G)$ by adding valid inequalities into the SDP formulation used for its computation. The computations are carried out on an Intel(R) Core(TM) i7-1260P CPU @ 2.10GHz with 64GB of RAM. We use Matlab for implementation, with MOSEK serving as the SDP solver.
4.1 The Stable Set Problem
We strengthen the Lovász theta function $\vartheta(G)$ towards the stability number $\alpha(G)$ by adding several classes of inequalities to the SDP for its computation (7). Specifically, we add the non-negativity constraints (10), the triangle inequalities (11) and (12), the inequalities valid for join of cliques (15), (16), and the newly introduced inequalities (18) and (19). The addition of inequalities is done in two phases.
  • Phase 1: In this phase, we incorporate well-known inequalities that have already been computationally investigated by other researchers. After computing the optimal solution of the Lovász theta function (7), we examine the obtained solution and search for violations of the non-negativity constraints (10) as well as the triangle inequalities (11) and (12). Since the number of found violations can be quite large, we add violated inequalities iteratively to the SDP (7). Specifically, we add all violations of the non-negativity constraints found in each iteration. For the triangle inequalities (11) and (12), we consider violations greater than $0.025$, and we add up to $2n$ of the largest violations per inequality type and per iteration. We iterate until there are fewer than $n$ violations of all considered inequalities, or until the number of iterations reaches $10$. The obtained upper bound on $\alpha(G)$ should satisfy almost all inequalities (10), (11), and (12). We denote this bound by BOUND 1.
  • Phase 2: In the second phase, we add inequalities (15) and (16) valid for the join of cliques as well as the newly proposed inequalities (18), (19) and (20). To separate these inequalities efficiently, we enumerate all cliques with at most five vertices and consider only chordless cycles of length $5$. The procedure for adding these inequalities is as follows: starting from the final solution of the SDP from Phase 1, we examine violations of the mentioned inequalities (15), (16), (18), (19) and (20). Then, similar to the procedure from Phase 1, we consider violations of these inequalities greater than $0.025$, and we add up to $2n$ of the largest violations per inequality type and per iteration. This procedure is repeated until there are fewer than $n$ violations of all considered inequalities, or until the number of iterations reaches $10$. We denote the resulting upper bound on $\alpha(G)$ by BOUND 2.
4.1.1 First Set of Experiments
In our first set of experiments, we deal with some instances considered by Gaar and Rendl in [12]. To our knowledge, the bounds presented in [12] are among the best SDP-based upper bounds on stability numbers available in the literature.
Specifically, we consider the following instances from [12]:
  • Near-regular graphs: These graphs were generated by Gaar and Rendl in [12] by the following procedure. First, they selected a perfect matching on $nr$ vertices. Next, they combined consecutive groups of $r$ vertices into a single vertex, which resulted in a regular multigraph on $n$ vertices. Finally, to obtain near-regular graphs, they removed loops and multiple edges from this multigraph.
  • Random graphs from the Erdős–Rényi model $G(n,p)$: In this model, the number of vertices $n$ is fixed, and each edge is included with probability $p$ independently of all other edges.
  • Torus graphs: These instances are constructed as follows. For a given $d$, the graph $T_{d}$ has $d^{2}$ vertices labeled by $(i,j)$ for $i$, $\in\{1,\ldots,d\}$. The vertical edges connect vertices with neighboring $i$ indices (keeping $j$ fixed), creating edges $\{(i,j),(i+1,j)\}$ modulo $d$. Similarly, the horizontal edges connect vertices with neighboring $j$ indices (keeping $i$ fixed), resulting in edges of the form $\{(i,j),(i,j+1)\}$ modulo $d$. Thus, the graph has $n=d^{2}$ vertices and $m=2n$ edges. As noted in [12], for even $d$, $\vartheta(T_{d})=\alpha(T_{d})$. Therefore, the study in [12] focused on torus graphs with odd $d$.
The goal of this computational study is to investigate the quality of bounds BOUND 1 and BOUND 2 as upper bounds on $\alpha(G)$ and to compare them to those presented in [12], which we denote by GR. Specifically, we focus on two key aspects: first, the improvement of BOUND 2 over BOUND 1, given that BOUND 2 is a refinement of BOUND 1; and second, the comparison between BOUND 2 and GR. Additionally, we assess the computational times required for these bounds. Although our computational setup differs from that in [12], we include the reported computational times from [12] to provide context for the computational efforts involved.
The results for the near-regular graphs are presented in Table 1. The first columns of Table 1 provide general information about the instances, including the name of the instance, the number of vertices $n$, the number of edges $m$, the value of $\alpha(G)$ as given in [12], and the value of the Lovász theta function $\vartheta(G)$. Next, we report the computed bounds BOUND 1 and BOUND 2, along with the bounds GR from [12]. For each instance, the best-computed bound on $\alpha(G)$ is highlighted in bold. Additionally, if BOUND 2 yields a better integer bound on $\alpha(G)$ than bound GR, the corresponding cell is shaded in gray. Finally, we present the computational times (time) required for the computation of bounds BOUND 1 and BOUND 2, as well as the times for the computation of the GR bound, as reported in [12].
Table 1: Stable set results for near-regular graphs considered in [12]
Graph $n$ $m$ $\alpha(G)\geq$ $\vartheta(G)$ BOUND 1 (time) BOUND 2 (time) GR (time) reg_n100_r4 100 195 40 43.449 41.246 (17) 40.713 (41) 40.687 (1164) reg_n100_r6 100 294 34 37.815 36.224 (7) 35.047 (32) 35.246 (1062) reg_n100_r8 100 377 31 34.480 33.337 (4) 32.063 (21) 32.190 (1067) reg_n200_r4 200 400 80 87.759 83.498 (111) 82.246 (260) 83.772 (1437) reg_n200_r6 200 593 68 79.276 76.047 (25) 73.709 (229) 75.555 (1523) reg_n200_r8 200 792 60 70.790 69.110 (11) 66.789 (78) 67.785 (1944) reg_n200_r10 200 980 57 66.418 65.142 (6) 62.695 (75) 62.894 (2556)
From the results for near-regular graphs presented in Table 1, it is evident that BOUND 2 provides significantly tighter estimates on $\alpha(G)$ compared to BOUND 1 across all tested instances. Notably, BOUND 2 outperforms BOUND 1 by at least one integer for every instance considered. Interestingly, despite this increased accuracy, BOUND 2 maintains a generally low computational cost.
Additionally, BOUND 2 frequently outperforms the GR bound, delivering better bounds in six out of seven cases. Even when BOUND 2 and the GR bound are equal, both bounds yield the same integer upper bound, which coincides with $\alpha(G)$. Furthermore, BOUND 2 is more computationally efficient than the GR method, making it a strong choice for improving upper bounds on the stability number of near-regular graphs.
Table 2: Stable set results for graphs from the Erdős–Rényi model considered in [12]
Graph $n$ $m$ $\alpha(G)$ $\vartheta(G)$ BOUND 1 (time) BOUND 2 (time) GR (time) rand_n100_p004 100 212 45 46.067 45.032 (22) 45.032 (1) 45.021 (432) rand_n100_p006 100 303 38 40.361 38.909 (13) 38.435 (20) 38.439 (887) rand_n100_p008 100 443 32 34.847 33.575 (5) 32.433 (23) 32.579 (1262) rand_n100_p010 100 489 32 34.020 32.934 (5) 32.151 (17) 32.191 (1138) rand_n200_p002 200 407 95 95.778 95.044 (222) 95.044 (1) 95.032 (836) rand_n200_p003 200 631 80 83.662 81.560 (39) 81.079 (52) 81.224 (1867) rand_n200_p004 200 816 67 73.908 71.654 (17) 69.818 (96) 70.839 (2227) rand_n200_p005 200 991 62 69.039 67.313 (19) 65.544 (70) 66.091 (2411)
The results for random graphs mirror those for near-regular graphs, with BOUND 2 consistently providing tighter estimates of $\alpha(G)$ compared to BOUND 1, as shown in Table 2. BOUND 2 also frequently outperforms the GR bound from [12], yielding better upper bounds on $\alpha(G)$ in six out of eight instances. Even in cases where GR performs better, both methods yield the same integer bound, matching $\alpha(G)$. Additionally, BOUND 2 offers a significant computational advantage, effectively delivering tighter upper bounds on the stability numbers of random graphs while reducing computational effort.
Table 3: Stable set results for torus graphs considered in [12]
Graph $n$ $m$ $\alpha(G)$ $\vartheta(G)$ BOUND 1 (time) BOUND 2 (time) GR (time) torus_5 25 50 10 11.180 10.000 (1) 10.000 (1) 10.002 (33) torus_7 49 98 21 23.224 21.000 (2) 21.000 (1) 21.009 (127) torus_9 81 162 36 39.241 36.000 (24) 36.000 (7) 36.021 (344) torus_11 121 242 55 59.249 55.022 (81) 55.019 (19) 55.066 (851) torus_13 169 338 78 83.254 78.379 (337) 78.048 (129) 79.084 (1031) torus_15 225 450 105 111.257 108.208 (1517) 105.214 (504) 106.287 (1615)
The results for two-dimensional torus graphs presented in Table 3 show that BOUND 2 consistently provides the tightest upper bounds on $\alpha(G)$ across all six instances, outperforming the GR bound and delivering integer bounds that exactly match $\alpha(G)$ for all evaluated torus graphs. Interestingly, BOUND 1 also proves to be highly competitive for torus graphs, often offering bounds that are better or very close to those of GR while requiring significantly less computational effort.
In [12], Gaar and Rendl emphasized the high computational demands of calculating upper bounds on the stability numbers of torus graphs, suggesting the need for a specialized SDP solver. Our findings indicate that incorporating non-negativity constraints and triangle inequalities into the SDP for the Lovász theta function already improves the precision of these upper bounds while reducing computational time. Therefore, any specialized SDP solver developed for these instances should integrate these inequalities to enhance both accuracy and efficiency.
4.1.2 Second Set of Experiments
In our second set of experiments, we investigate several well-established instances from the literature:
  • DIMACS Instances: We analyze several benchmarks from the Second DIMACS Implementation Challenge [22], held in 1992 and 1993, which focused on NP-hard problems including maximum clique, graph coloring, and satisfiability. For our study on the stable set problem, we use the complements of the maximum clique instances from this challenge.
  • Spin graphs: The stability numbers of certain spin graphs were recently analyzed using an SDP-based branch and bound framework in a work by Gaar, Siebenhofer, and Wiegele [13]. These instances are particularly interesting because, for most of the spin graphs studied in [13], only lower and upper bounds on $\alpha(G)$ were provided. Therefore, it is valuable to investigate whether the upper bounds on $\alpha(G)$ can be improved by strengthening the Lovász theta function $\vartheta(G)$ through the addition of newly proposed inequalities and those valid for the join of cliques.
  • Evil graphs: We consider extremely hard and versatile (evil) instances for the maximum clique problem recently introduced by Szabo and Zaválnij in [40], known for their wide range of difficulties. These instances, comprising a total of $40$ graphs, are highly valued in algorithmic research. The clique numbers of evil graphs were investigated by San Segundo, Furini, Álvarez, and Pardalos in [37]. Since our focus is on the stable set problem, we consider the complements of some of these instances. To our knowledge, this is the first computational study of the SDP-based upper bounds on the stability numbers of evil instances.
The primary objective of this computational study is to evaluate the extent to which BOUND 2 improves upon BOUND 1 as an upper bound on $\alpha(G)$. Additionally, we examine the computational times required to calculate both BOUND 1 and BOUND 2.
The computational results for the spin graphs and selected DIMACS instances are presented in Table 4. The table follows the same structure as those from the first set of experiments, with the exception that the GR bound is not reported. Additionally, we know that BOUND 2 outperforms BOUND 1 as an upper bound on $\alpha(G)$ and that is why we do not use bold formatting to highlight the best computed bound for each instance. However, if BOUND 2 surpasses BOUND 1 by at least one integer value, we highlight this result in gray.
Table 4: Stable set results for selected DIMACS instances and spin graphs
Graph $n$ $m$ $\alpha(G)$ $\vartheta(G)$ BOUND 1 (time) BOUND 2 (time) spin5 125 375 50 55.902 50.000 (17) 50.000 (6) spin7 343 1029 $147\leq\alpha\leq 151$ 162.566 147.000 (1225) 147.000 (639) MANN_a9 45 72 16 17.475 17.220 (2) 17.220 (1) MANN_a27 378 702 126 132.763 131.709 (781) 131.112 (874) C125.9 125 787 34 37.805 36.920 (4) 35.568 (32) C250.9 250 3141 44 56.241 55.771 (18) 54.899 (479) sanr200_0_9 200 2037 42 49.274 48.723 (11) 47.472 (229)
From the results for spin graphs presented in Table 4, we observe that the values of BOUND 1 and BOUND 2 coincide with the stability numbers of considered graphs. This is particularly noteworthy as it indicates that we found the optimal stability number for the spin7 graph. Previously, only lower and upper bounds for $\alpha(G)$ were reported in [13]. These results highlight the potential for further research into SDP-based upper bounds for spin graphs, especially through the enhancement of the SDP by incorporating additional valid inequalities to refine the computation of the Lovász theta function (7).
For the DIMACS instances, as shown in Table 4, BOUND 1 already significantly improves upon the Lovász theta function as an upper bound on $\alpha(G)$ for most of the considered graphs. BOUND 2 further enhances these bounds, offering notable improvements over BOUND 1 for several instances. In particular, BOUND 2 provides better integer bounds on $\alpha(G)$ for three of the graphs. Additionally, the computations for both BOUND 1 and BOUND 2 were executed efficiently, with the only exception being the graph MANN_a27.
Table 5: Stable set results for selected evil instances
Graph $n$ $m$ $\alpha(G)$ $\vartheta(G)$ BOUND 1 (time) BOUND 2 (time) evil-N120-p98-chv12x10 120 545 20 24.526 24.526 (1) 20.000 (2) evil-N120-p98-myc5x24 120 236 48 52.607 48.000 (22) 48.000 (16) evil-N121-p98-myc11x11 121 508 22 26.397 26.397 (1) 22.000 (1) evil-N125-p98-s3m25x5 125 873 20 25.000 22.361 (1) 22.361 (5) evil-N138-p98-myc23x6 138 1242 12 15.177 15.177 (1) 15.177 (3) evil-N150-p98-myc5x30 150 338 60 65.121 60.000 (32) 60.000 (43) evil-N150-p98-s3m25x6 150 1102 24 30.000 26.833 (2) 26.833 (8) evil-N154-p98-myc11x14 154 701 28 33.596 33.596 (1) 28.000 (2) evil-N180-p98-chv12x15 180 944 30 36.788 36.788 (1) 30.000 (5) evil-N184-p98-myc23x8 184 1764 16 20.235 20.235 (2) 20.235 (7)
Finally, we examine the results for evil instances presented in Table 5. As this is the first study of SDP-based bounds for these instances, we also review the values of the Lovász theta function, $\vartheta(G)$. Notably, $\vartheta(G)$ consistently overestimates $\alpha(G)$ across all instances, with differences ranging from approximately $3$ to $6.8$, indicating its lack of tightness as an upper bound. BOUND 1 provided a tighter upper bound on $\alpha(G)$ than $\vartheta(G)$ in four out of ten instances, and in some cases, it exactly matched $\alpha(G)$. BOUND 2 demonstrates significant improvements over BOUND 1. Notably, for six out of the ten instances considered, BOUND 2 exactly matches the stability number $\alpha(G)$.
Interestingly, in instances where BOUND 2 does not exactly match $\alpha(G)$, it provides the same value as BOUND 1. In particular, we observe that for instances evil-N138-p98-myc23x6 and evil-N184-p98-myc23x8 both BOUND 1 and BOUND 2 coincide with $\vartheta(G)$. Therefore, it is a future research to explore whether incorporating some other inequalities into the SDP to compute the Lovász theta function might improve the bounds on stability numbers for these instances.
Finally, we note that the computational effort required to calculate BOUND 1 and BOUND 2 was relatively modest. In particular, we obtained stability numbers for several instances in remarkably short running times for most instances. This not only highlights the quality of the results but also emphasizes the rapid computation of upper bounds on the stability numbers of evil graphs.
4.2 The Graph Coloring Problem
We conduct a computational study to strengthen the Lovász theta function $\vartheta(\overline{G})$ towards $\chi(G)$ in the same manner as we did for the stable set problem in Section 4.1. Hence, we iteratively add inequalities (13), (14), (17), (21), and (23) into the SDP for computing the Lovász theta function (9) in two phases.
  • Phase 1: Starting from the optimal solution of the SDP (9), in each iteration, we add all violations of the non-negativity constraints (13) and up to $2n$ of the largest violations greater than $0.025$ of the triangle inequalities (14). This process continues until there are fewer than $n$ violations or until the number of iterations reaches $10$. The resulting lower bound on $\chi(G)$ is denoted by BOUND 1.
  • Phase 2: Using the optimal solution from the SDP to compute BOUND 1, we then address violations of inequalities valid for a clique and one additional vertex (17) as well as the newly proposed inequalities (21) and (23). For this purpose, we enumerate all cliques with at most $5$ vertices and consider induced cycles of length $5$. In each iteration, we add up to $2n$ of the largest violations greater than $0.025$ per inequality type. This phase continues until there are fewer than $n$ violations or until the number of iterations reaches 10. We denote the resulting lower bound on $\chi(G)$ by BOUND 2.
4.2.1 First Set of Experiments
For the first set of experiments, we consider some of the instances analyzed by Gaar and Rendl in [12]. Specifically, we examine a mug graph on 88 vertices, several Mycielski graphs, and several Full Insertion graphs, which are a generalization of Mycielski graphs. Mycielski graphs, introduced by Tomescu in [42], are particularly interesting because they are challenging to solve. Since they are triangle-free, their clique number is $2$, but their chromatic number increases with problem size.
The primary objective of this computational study is to evaluate the improvement of BOUND 2 over BOUND 1 as a lower bound on $\chi(G)$ and to compare our results with those obtained in [12]. Additionally, we aim to investigate the computational times required to compute these bounds.
The results for this first set of experiments are given in Table 5. The presentation of these results follows the same format as the data for the stable set problem in Section 4.1.
Table 6: Graph coloring results for some instances considered in [12]
Graph $n$ $m$ $\chi(G)$ $\vartheta(\overline{G})$ BOUND 1 (time) BOUND 2 (time) GR (time) myciel5 47 236 6 2.639 3.093 (3) 3.468 (17) 3.510 (4240) myciel6 95 755 7 2.734 3.253 (21) 3.622 (406) 3.534 (1540) mug88_1 88 146 4 3.000 3.001 (3) 3.001 (1) 3.022 (4709) 1_FullIns_4 93 593 5 3.124 3.487 (3) 3.837 (23) 3.939 (7220) 2_FullIns_4 212 1621 6 4.056 4.343 (4) 4.670 (17) 4.700 (10106)
The results in Table 6 show that BOUND 1 consistently improves upon $\vartheta(\overline{G})$ as a lower bound on $\chi(G)$ across all five instances, with BOUND 2 offering additional enhancements over BOUND 1 in four out of five cases. When comparing BOUND 2 as a lower bound on $\chi(G)$ to the GR bounds, BOUND 2 often comes very close to the GR results and for one instance, BOUND 2 actually surpasses GR. Nevertheless, for all considered instances, BOUND 2 provides the same integer bounds on $\chi(G)$ as the GR bound.
An important consideration is the computational time required for each method. The GR bounds, while often slightly superior, come at a significantly higher computational cost. This efficiency underscores practical benefit of BOUND 2, providing competitive bounds with far less computational effort. Future research should investigate further inequalities to potentially enhance the performance of BOUND 2 and achieve even better results within shorter computation times.
4.2.2 Second Set of Experimens
In our second set of experiments, we consider some instances from the literature:
  • DSJC graphs: DSJC graphs are standard $G(n,p)$ random graphs. They were initially used in the paper by Johnson, Aragon, McGeoch, and Schevon in [21] and were also included in the second DIMACS challenge. We consider two sparse DSJC graphs.
  • Full Insertion graphs: We explore additional Full Insertion graphs that were not included in [12].
  • Queen graphs: The $n\times n$ queen graph is a graph on $n^{2}$ vertices in which each vertex represents a square in an $n\times n$ chessboard, and each edge corresponds to a legal move by a queen.
  • Random graphs from [10]: We examine some of the random graphs considered by Dukanovic and Rendl in [10]. Since they strengthened $\overline{\vartheta}(G)$ towards $\chi(G)$ by adding the non-negativity constraints as well as triangle inequalities, their bounds should generally coincide with our BOUND 1, so we do not report their results separately. Moreover, we do not know the exact values of $\chi(G)$ for these instances, hence these values are not reported.
In this computational study, we compare the values of $\vartheta(\overline{G})$, BOUND 1, and BOUND 2 as lower bounds on $\chi(G)$ and assess the computational effort required for the computations of BOUND 1 and BOUND 2. The results are presented in Table 7.
Table 7: Graph coloring results for selected instances from the literature
Graph $n$ $m$ $\chi(G)$ $\vartheta(\overline{G})$ BOUND 1 (time) BOUND 2 (time) dsjc125.1 125 736 5 4.106 4.218 (2) 4.430 (14) dsjc250.1 250 3218 8 4.906 4.939 (12) 5.040 (457) 3_FullIns_3 80 346 6 5.016 5.194 (1) 5.194 (1) 4_FullIns_3 114 541 7 6.010 6.010 (1) 6.309 (1) 5_FullIns_3 154 792 8 7.007 7.007 (1) 7.267 (2) Queen_8_8 64 728 9 8.000 8.000 (1) 8.000 (7) Queen_9_9 81 1056 10 9.000 9.000 (1) 9.000 (25) Queen_10_10 100 1470 11 10.000 10.000 (1) 10.000 (62) G100_25 100 1240 - 5.823 5.867 (2) 6.235 (76) G150_25 150 2802 - 6.864 6.918 (6) 7.184 (529) G200_1 200 2047 - 4.447 4.473 (10) 4.600 (178) G250_1 250 3149 - 4.805 4.831 (12) 4.928 (512)
The results in Table 7 indicate that BOUND 1 improves upon $\vartheta(\overline{G})$ as a lower bound on $\chi(G)$ for seven out of twelve instances. BOUND 2, in comparison, provides better results for eight instances and surpasses BOUND 1 by one integer value in three cases, demonstrating its effectiveness in tightening lower bounds.
However, BOUND 2 requires more computational effort, especially for instances where it improves upon BOUND 1 by one integer value. Notably, for all queen graphs considered, both bounds align with the value of the Lovász theta function $\vartheta(\overline{G})$, suggesting that the added inequalities do not enhance this bound. Further research could explore whether other inequalities might improve $\vartheta(\overline{G})$ as a lower bound on $\chi(G)$.
4.3 Summary of Results
In Subsections 4.1 and 4.2, we performed an extensive computational study of bounds for stability and chromatic numbers by incorporating valid inequalities into the SDP for the computation of the Lovász theta function. We assessed these bounds across various instances from the literature and provided detailed analysis.
Overall, our findings show that once BOUND 1 is computed by incorporating non-negativity constraints and triangle inequalities into the SDP for the computation of the Lovász theta function, computing BOUND 2 becomes both efficient and straightforward. Furthermore, BOUND 2 not only computes quickly and easily but also significantly improves upon BOUND 1 for several instance classes, often providing better integer solutions, as seen particularly in Subsection 4.1. Notably, we limited our iterative computational approach to a maximum of ten iterations, but only a few iterations were required for most instances, further demonstrating the practical value of this approach for deriving upper bounds on $\alpha(G)$ and lower bounds on $\chi(G)$.
We observe that the effectiveness of the computed bounds on $\alpha(G)$ and $\chi(G)$ generally depends on the size and sparsity of the instance. The best results are obtained for smaller and sparser instances, while performance tends to decrease for larger and denser ones. Interestingly, for certain evil instances in Subsection 4.1 as well as queen graphs in Subsection 4.2, adding inequalities into the SDP did not enhance the Lovász theta function. Despite this, our bounds remain highly competitive with those found in the literature and can be computed in relatively short running times.
5 Conclusion
Several questions remain open for future investigation. First, it would be valuable to explore which additional inequalities could be added into the SDP for the computation of the Lovász theta function to improve upper bounds on $\alpha(G)$ for certain evil graphs and lower bounds on $\chi(G)$ for queen graphs. Additionally, our research thus far has concentrated on subgraphs containing cliques and odd cycles. It is an open question to identify other easily recognizable structures and examine whether incorporating valid inequalities for these structures could further enhance the Lovász theta function with respect to $\alpha(G)$ and $\chi(G)$. Regarding the new inequalities introduced in this work, it would be beneficial to determine which additional inequalities are needed to fully characterize the underlying polytopes, as we have done for subgraphs containing cliques in [35]. Finally, given the strength of the proposed bounds, developing an exact algorithm for the stable set problem within a branch-and-bound framework that utilizes these bounds would be of significant interest.
Disclosure statement
The authors report there are no competing interests to declare.
Data availability statement
The program code associated with this paper is available as ancillary files from the arXiv page of this paper. Additionally, the source code and data are available upon request from the authors.
Funding
This research was funded in whole by the Austrian Science Fund (FWF) [10.55776/DOC78]. For the purposes of open access, the authors have applied a CC BY public copyright license to all author-accepted manuscript versions resulting from this submission.